OpenStudy (anonymous):

solve the differential equation (i) (x+y)^2 dy/dx=a^2 (put x+y=z)

OpenStudy (anonymous):

expand, and let z = y/x

OpenStudy (anonymous):

okay can u do it for me to understand it very well

OpenStudy (anonymous):

(x+y)^2 = x^2 + y^2 + 2xy divide both sides by x^2 [1 + (y/x)^2 + 2(y/x)] (dy/dx) = a^2/x^2 z = y/x,

OpenStudy (anonymous):

okay then what i have to do ??